x-squared + 10x - 7 = 0 , completing the square
Answer by ad_alta(240)

1) x^2+10x-7=0
2) [x+(5)]^2=7+25
3) x+5= [+/-] (32)^(1/2)
4)x= -5 [+/-] 4*sqrt(2)